LeBeanMan:

Given the functions a(x) = 3x − 12 and b(x) = x − 9, solve a[b(x)]. a[b(x)] = 3x2 − 21 a[b(x)] = 3x2 − 39 a[b(x)] = 3x − 21 a[b(x)] = 3x − 39

xxDeppressionxx:

a(x) = 3x - 12 b(x) = x - 9 a(b(x)) = a(x - 9) = 3(x - 9) - 12 = 3x - 27 - 12 = 3x - 39 a(b(x)) = 3x - 39 option (d) is correct: (a[b(x)] = 3x − 39)

LeBeanMan:

Thx :)

xxDeppressionxx:

Np